The coin is dropped by a man standing in the lift, 1)When the lift is at rest, then it hit the floor in T1sec 2)When the lift is moving upward with acceleration =a,then it hit the floor in T2sec 3)When the lift is moving downward with acceleration =a,then it hit the floor in T3sec, then find out the relation between T1,T2 and T3.

Grade:11

1 Answers

Vikas TU
5 years ago
  1. When the lift is at rest.
h = 0 + 0.5gt1^2
ย 
  1. When the lift is accelearted by a upwards,
h = 0 + 0.5(g+a)t2^2
ย 
  1. When the lift is accelerated downwards,
h = 0 + 0.5(g โ€“ a)t3^2
ย 
From all the 3 equations if we relate we get,
(g-a)t3^2 = (g+a)t2^2
or
(t2/t3)^2 = (g-a)/(g+a)
t2/t3 = root(g-a/g+a)
ย 
For t1,
ย 0.5(g+a)t2^2 =ย 0.5g(t1^2)
t1/t2 = root((g+a)/g)
ย 
Hence these would all be the relations.